# ThinK: Thinner Key Cache by Query-Driven Pruning

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications. However, their increased computational and memory demands present significant challenges, especially when handling long sequences. This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference. Unlike existing approaches that optimize the memory based on the sequence length, we identify substantial redundancy in the channel dimension of the KV cache, as indicated by an uneven magnitude distribution and a low-rank structure in the attention weights. In response, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels. Our approach not only maintains or enhances model accuracy but also achieves a reduction in KV cache memory costs by over 20% compared with vanilla KV cache eviction and quantization methods. For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU. Extensive evaluations on the LLaMA and Mistral models across various long-sequence datasets verified the efficiency of ThinK, establishing a new baseline algorithm for efficient LLM deployment without compromising performance.
@@ 1 Introduction
Despite their success in various applications, generating outputs with LLMs incurs significant computational and financial costs, which rise with increasing model size and sequence length. Both the training (Strubell et al., 2020; Hoffmann et al., 2022; Dong et al., 2024a) and inference (Ainslie et al., 2023) stages involve frequent generation, further contributing to these costs. Consequently, efficient LLMs have gained traction in recent years (Hu et al., 2021; Wan et al., 2023). To address these challenges, quantization (Frantar et al., 2022; Lin et al., 2024; Dettmers et al., 2024; Xu et al., 2023) and pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023) are employed to reduce model size. Additionally, the key-value (KV) cache, stored in GPU memory alongside model parameters, scales linearly with both sequence length and batch size, creating a substantial memory burden when handling long sequences. Consequently, effective management of extended contexts is essential for the practical deployment of LLMs. In this paper, we focus on the long-context scenario, aiming to reduce memory consumption associated with processing lengthy sequences.
@@ 3.1 Preliminary Study of KV Cache Optimization
In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size. Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows: $2\times B\times S\times L\times N\times D$, where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension. The KV cach size grows linearly as the batch size $B$ and sequence length $S$. For a model with multihead attention (MHA) (Vaswani et al., 2017), such as LLaMA2-7B (Touvron et al., 2023b), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters. The KV cache must be transferred from off-chip memory (HBM) (Jia et al., 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck. This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies. Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b; Zhang et al., 2024c; Li et al., 2024) and precision (Hooper et al., 2024; Liu et al., 2024b). We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
@@
In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA-3-8B model. We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al., 2023). Compared to the baseline methods, H2O (Zhang et al., 2024c) and SnapKV (Li et al., 2024), both with a KV length of 512, we further prune the channels of the key cache. A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning. The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension. These results also indicate the need for a better pruning matrix to achieve higher pruning ratios effectively.
@@ 3.2 Query-Driven Pruning
Observation Window. Based on the observations in SnapKV (Li et al., 2024) that the last window of input sequence recognizes highly similar attention pattern with generation. To reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
@@ 3.3 Implementation of ThinK
Following SnapKV, we focus on the long context input scenario. We opt not to prune the most recent tokens and newly generated keys. Consequently, our key-value (KV) cache will store two distinct categories of keys: one subset consists of pruned keys with a reduced channel size, while the other retains keys at their original size. Additionally, we maintain a binary mask to indicate which channels have been pruned. it is important to note that the memory overhead associated with this mask is negligible. Figure 2 illustrates one implementation of ThinK during the decoding stage. This implementation begins by pruning the query using the mask. The pruned query is then multiplied by the pruned key, while the unpruned query is applied to the unpruned Key. Subsequently, the two outputs are concatenated.
@@ 4 Experiment Results
In this section, we conduct comprehensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction. First, we introduce benchmark datasets, baseline methods, and implementation details in Section 4.1. Next, we report the performance ThinK with two KV eviction methods and one KV quantization methods in long-context scenarios in Section 4.2 and  4.3. Finally, we present the results of ablation experiments in Section 4.4.
@@ 4.1 Settings
Implementation Details. In this paper, we use LLaMA-2-7B-chat, LLaMA-3-8B-Instruct, LLaMA-3-70B-Instruct (Meta, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al., 2020). Our ThinK aims to prune channels of the key cache, which is agnostic to KV cache compression methods. If there is no other statement, we prune the key cache by default. All the experiments are conducted using NVIDIA A100 GPUs. To ensure a fair comparison between KV cache compression methods and their integration with ThinK, we applied consistent hyperparameters across both settings. For instance, when comparing SnapKV and SnapKV integrated with ThinK, we used a maximum pooling kernel size of 7 and an observation window size of 32, maintaining the same KV-size for both configurations.
@@ 4.2 Results on LongBench
Tables 2 and  3 present the results of KV compression methods and their integration with our proposed channel pruning technique for the key cache (ThinK) across three different base LLMs, evaluated at various KV-sizes on the LongBench benchmark. The following observations can be drawn: (1) Our method successfully prunes the channels of the key cache after the KV cache has been compressed using H2O and SnapKV. For the LLaMA-3-8B-Instruct base model, our approach reduces memory usage while slightly improving performance for both H2O and SnapKV. For the Mistral-7B-Instruct-v0.2 base model, our method similarly reduces memory usage, with only a minor performance drop in some cases. For the larger LLaMA-3-70B base model, our method achieves comparable or superior performance after pruning $40\%$ of the key cache channels, compared to the SnapKV baselines. (2) When Comparing SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm in Table 1, our query-driven channel pruning approach demonstrates superior performance when the pruning ratio of $\lambda=0.4$. (3) Lower pruning ratios generally result in better performance compared to higher pruning ratios. (4) As the KV-size increases from $128$ to $2048$, the performance of our channel pruning method improves. Notably, with a KV-size of $2048$ and a pruning ratio of $0.4$, our method even surpasses the performance of LLaMA-3-8B-Instruct with a full KV cache. These findings suggest that our method is agnostic to the underlying KV cache compression techniques and can further enhance both performance and memory efficiency. Moreover, query-driven channel pruning proves more effective than $l_{1}$ and $l_{2}$ norm-based methods for channel pruning in LLMs.
@@ 4.3 Results on Needle-in-a-Haystack
Figure 6 (a)-(d) (in Appendix B) visualize the Needle-in-a-Haystack test accuracy across varying token lengths and depth percentages. The KV-sizes are set to $128$ and $1024$, with pruning ratios of $\lambda=0.4$ and $\lambda=0.5$, respectively. ThinK preserves the retrieval capabilities of SnapKV, although there are minor numerical differences in overall accuracy (e.g., $77.8$ vs. $78.6$ and $90.4$ vs. $90.6$). ThinK matches SnapKV in accuracy for the majority of token limits and depths, demonstrating consistency in performance. Furthermore, ThinK successfully retrieves certain ”needles” that SnapKV fails to capture, resulting in improved overall accuracy. These visualizations highlight the robustness of ThinK from a fine-grained perspective, illustrating its capacity to enhance the original approach.
@@ 4.4 Ablation Studies
Memory Usage Comparison. To evaluate the efficiency of ThinK, we follow the methodology used in KIVI (Liu et al., 2024b). We generate synthetic workloads with an input prompt length of 160 and an output length of 338. The peak memory usage is reported for the vanilla FP16 baseline, KIVI, and KIVI combined with ThinK ($0.4$) for LLaMA-2-7B-chat. As in Figure 3(b), the memory savings from our method become increasingly evident as the batch size grows, in both the KIVI $/2$ 2 and KIVI $/4$ 4 configurations. Compared to the baseline model, our approach achieves over a 5$\times$ increase in batch size while maintaining the same memory footprint when integrated with KIVI.
Pruning Channels of Both Key and Value Cache. In our previous study, we evaluate methods that focused on pruning channels of the key cache. In this part, we explore the impact of pruning channels in the value cache as well (Appendix D). Specifically, for KV cache compression methods, we apply different pruning ratios to the channels of the key and value caches. Table 8 presents the results with LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2 on the LongBench benchmark. When evaluating the base model LLaMA-3-8B-Instruct, H2O or SnapKV with both key and value channel pruning perform comparably to H2O or SnapKV without pruning. In certain instances, the models with key and value channel pruning even outperform their non-pruned counterparts. For the base model Mistral-7B-Instruct-v0.2, pruning the value cache channels leads to a slight performance drop. Nevertheless, it is important to note that applying value channel pruning to the KV cache further reduces memory usage.
@@ 5 Related Work
KV Cache Eviction. StreamingLLM (Xiao et al., 2023b) retains a few initial tokens along with some recent tokens based on the observation of attention sinks, which can leading to the loss of critical information carried by the dropped tokens. H2O (Zhang et al., 2024c) selectively retains a small subset of tokens by greedily dropping those with lower contributions to cumulative attention. SnapKV (Li et al., 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts. FastGen (Ge et al., 2023) adaptively evicts tokens from attention heads that focus on local contexts, discarding non-special tokens that surround key tokens, while standard KV cache is applied to attention heads that attend more broadly. PyramidKV (Zhang et al., 2024b) and PyramidInfer (Yang et al., 2024) take a hierarchical approach, adjusting KV cache sizes across different layers by allocating more cache to lower layers and less to higher ones.
@@
Structured Pruning of LLMs. Traditional structured pruning  (Ma et al., 2023; Ding et al., 2023) of LLMs typically focuses on removing unimportant layers, heads, or hidden dimensions, often leading to significant performance degradation. In contrast, our approach preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache. By dynamically identifying unimportant channels using data dependant criterion, our method greatly reduce the key cache size with minimal performance loss.
6 Conclusion
Inspired by the observation that certain channels exhibit significantly larger magnitudes than others, and supported by singular value analysis indicating that the key cache is inherently low-rank, we propose ThinK  as a pruning method targeting the key cache channels. Our pruning strategy is query-dependent and optimized based on attention scores, ensuring that essential information is preserved for each input query. In addition, ThinK can be seamlessly integrated with existing token-level KV cache pruning techniques (Li et al., 2024; Zhang et al., 2024c) and KV cache quantization methods (Liu et al., 2024b), further enhancing inference efficiency. Extensive experiments on LongBench and Needle-in-a-Haystack benchmarks demonstrate the effectiveness our query-dependent channel pruning approach. Our method achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%. For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU.

### v3

@@ Abstract
Large Language Models (LLMs) have revolutionized the field of natural language processing, achieving unprecedented performance across a variety of applications. However, their increased computational and memory demands present significant challenges, especially when handling long sequences. This paper focuses on the long-context scenario, addressing the inefficiencies in KV cache memory consumption during inference. Unlike existing approaches that optimize the memory based on the sequence length, we identify substantial redundancy in the channel dimension of the KV cache, as indicated by an uneven magnitude distribution and a low-rank structure in the attention weights. In response, we propose ThinK, a novel query-dependent KV cache pruning method designed to minimize attention weight loss while selectively pruning the least significant channels. Our approach not only maintains or enhances model accuracy but also achieves a reduction in KV cache memory costs by over 20% compared with vanilla KV cache eviction and quantization methods. For instance, ThinK integrated with KIVI can achieve 2.8×2.8$\times$2.8 × peak memory reduction while maintaining nearly the same quality, enabling a batch size increase from 4$\times$ (with KIVI alone) to 5$\times$ when using a single GPU. Extensive evaluations on the LLaMA and Mistral models across various long-sequence datasets verified the efficiency of ThinK. Our code has been made available at https://github.com/SalesforceAIResearch/ThinK.
@@ 1 Introduction
Despite their success in various applications, generating outputs with LLMs incurs significant computational and financial costs, which rise with increasing model size and sequence length. Both the training (Strubell et al., 2020; Hoffmann et al., 2022; Dong et al., 2024a) and inference (Ainslie et al., 2023) stages involve frequent generation, further contributing to these costs. Consequently, efficient LLMs have gained traction in recent years (Hu et al., 2021; Wan et al., 2023). To address these challenges, quantization (Frantar et al., 2022; Lin et al., 2024; Dettmers et al., 2024; Xu et al., 2023) and pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023; Lu et al., 2024b) are employed to reduce model size. Additionally, the key-value (KV) cache, stored in GPU memory alongside model parameters, scales linearly with both sequence length and batch size, creating a substantial memory burden when handling long sequences. Consequently, effective management of extended contexts is essential for the practical deployment of LLMs. In this paper, we focus on the long-context scenario, aiming to reduce memory consumption associated with processing lengthy sequences.
@@ 3.1 Preliminary Study of KV Cache Optimization
In scenarios with extended contexts or batch processing, the main limitations in terms of memory and speed are due to the handling of the KV cache size. Considering a batch of requests to a Large Language Model (LLM) service that provides a long input prompt consisting of tokens $[x_{B1},...,x_{BS}]$, the total KV cache size can be computed as follows: $2\times B\times S\times L\times N\times D$, where $L$ is the number of layers, $N$ is the number of heads, $D$ is the head dimension. The KV cache size grows linearly as the batch size $B$ and sequence length $S$. For a model with multihead attention (MHA) (Vaswani et al., 2017), such as LLaMA2-7B (Touvron et al., 2023b), a context length of $2048$ and a batch size of $13$ require storing a $13$ GB KV cache, which is equivalent to the size of the model parameters. The KV cache must be transferred from off-chip memory (HBM) (Jia et al., 2018) to on-chip memory (cache) for each token generated, leading to a memory bottleneck. This substantial memory demand highlights the challenges in managing large-scale models and the need for efficient memory utilization strategies. Current methods optimize the KV cache based on the sequence length $S$ (Xiao et al., 2023b; Zhang et al., 2024c; Li et al., 2024) and precision (Hooper et al., 2024; Liu et al., 2024b). We will introduce a new method, ThinK, to optimize it from the perspective of the number of head dimensions $D$.
@@
In Table 1, we present the results of key cache pruning with various pruning ratios applied to the LLaMA-3-8B model. We utilize the $l_{1}$ and $l_{2}$ norms as criteria for evaluation, and validate performance using the LongBench benchmark (Bai et al., 2023). Compared to the baseline methods, H2O (Zhang et al., 2024c) and SnapKV (Li et al., 2024), both with a KV length of 512, we further prune the channels of the key cache. A $30\%$ pruning ratio can maintain accuracy; however, increasing it to $40\%$ results in significant performance degradation, especially for $l_{1}$ norm based pruning. The results of magnitude-based pruning support our assumption that the key cache is redundant in the channel dimension. These results also indicate the need for a better pruning metrics to achieve higher pruning ratios effectively.
@@ 3.2 Query-Driven Pruning
Observation Window. Following SnapKV (Li et al., 2024), to reduce the computation cost, we only use the last $S_{obs}$ window to calculate the score as the last window of input sequence recognizes highly similar attention pattern with generation: $\|\textbf{Q}_{i}[-S_{\text{obs}}:,j]\textbf{K}_{i}[:,j]^{T}\|_{F}$.
@@ 3.3 Implementation of ThinK
We opt not to prune the most recent tokens and newly generated keys. Consequently, our key-value (KV) cache has two categories: one subset consists of pruned keys with a reduced channel size, while the other at their original size. Additionally, we maintain a binary mask whose memory overhead is negligible to indicate which channels have been pruned. Figure 2 illustrates one implementation of ThinK during the decoding stage. We first prune the query using the stored mask, ensuring that the query dimensionality and the pruned key cache remain consistent. The pruned query is then multiplied by the pruned key, while the unpruned query is applied to the unpruned key. Subsequently, the two outputs are concatenated. Our method removes unimportant Key cache channels to get real memory reduction.
@@ 4 Experiment Results
In this section, we conduct comprehensive experiments to evaluate the effectiveness of ThinK on performance and memory reduction. In the experiments, we prune Key cache channels by default, the value cache results are in Table 8 in Appendix D.
@@ 4.1 Settings
Implementation Details. In this paper, we use LLaMA-2-7B-chat, LLaMA-3-8B-Instruct, LLaMA-3-70B-Instruct (Meta, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as the backbone LLMs, both accessible via HuggingFace (Wolf et al., 2020). Our ThinK aims to prune channels of the key cache, which is agnostic to KV cache compression methods. If there is no other statement, we prune the key cache by default. All the experiments are conducted using NVIDIA A100 GPUs. To ensure a fair comparison between KV cache compression methods and their integration with ThinK, we applied consistent hyperparameters across both settings. For instance, when comparing SnapKV and SnapKV integrated with ThinK, we used a maximum pooling kernel size of 7 and an observation window size of 32, maintaining the same KV-size for both configurations. We compress the Key cache starting from the prefilling stage.
@@ 4.2 Results on LongBench
Tables 2 and  3 present the results of KV compression methods and their integration with our proposed channel pruning technique for the key cache (ThinK) across three different base LLMs, evaluated at various KV-sizes on the LongBench benchmark. The pruning ratio of $\lambda=0.4$ indicate that 40% of key cache channels are removed ,resulting in a 20% reduction in the total KV cache memory footprint. The following observations can be drawn: (1) Our method successfully prunes the channels of the key cache after the KV cache has been compressed using H2O and SnapKV. For the LLaMA-3-8B-Instruct base model, our approach reduces memory usage while slightly improving performance for both H2O and SnapKV. For the Mistral-7B-Instruct-v0.2 base model, our method similarly reduces memory usage, with only a minor performance drop in some cases. For the larger LLaMA-3-70B base model, our method achieves comparable or superior performance after pruning $40\%$ of the key cache channels, compared to the SnapKV baselines. (2) When Comparing SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm in Table 1, our query-driven channel pruning approach demonstrates superior performance when the pruning ratio of $\lambda=0.4$. (3) Lower pruning ratios generally result in better performance compared to higher pruning ratios. (4) As the KV-size increases from $128$ to $2048$, the performance of our channel pruning method improves. Notably, with a KV-size of $2048$ and a pruning ratio of $0.4$, our method even surpasses the performance of LLaMA-3-8B-Instruct with a full KV cache. These findings suggest that our method is agnostic to the underlying KV cache compression techniques and can further enhance both performance and memory efficiency. Moreover, query-driven channel pruning proves more effective than $l_{1}$ and $l_{2}$ norm-based methods for channel pruning in LLMs. Experiment results of applying ThinK directly to vanilla models are in Table 9 which further validate the effectiveness of our method.
@@ 4.3 Results on Needle-in-a-Haystack
Figure 6 (a)-(d) (in Appendix B) visualize the Needle-in-a-Haystack test accuracy across varying token lengths and depths. The KV-sizes are set to $128$ and $1024$, with pruning ratios of $\lambda=0.4$ and $\lambda=0.5$, respectively. ThinK preserves the retrieval capabilities of SnapKV, although there are minor numerical differences in overall accuracy (e.g., $77.8$ vs. $78.6$ and $90.4$ vs. $90.6$). ThinK matches SnapKV in accuracy for the majority of token limits and depths, demonstrating consistency in performance. Furthermore, ThinK successfully retrieves certain ”needles” that SnapKV fails to capture, resulting in improved overall accuracy. These visualizations highlight the robustness of ThinK from a fine-grained perspective, illustrating its capacity to enhance the original approach.
@@ 4.4 Ablation Studies
Memory Usage Comparison. To evaluate the efficiency of ThinK, we follow the methodology used in KIVI (Liu et al., 2024b). We generate synthetic workloads with an input prompt length of 160 and an output length of 338. The peak memory usage is reported for the vanilla FP16 baseline, KIVI, and KIVI combined with ThinK ($0.4$) for LLaMA-2-7B-chat. As in Figure 3(b), the memory savings from our method become increasingly evident as the batch size grows, in both the KIVI $/2$ 2 and KIVI $/4$ 4 configurations. Compared to the baseline model, our approach achieves over a 5$\times$ (from 4$\times$ with KIVI alone) increase in batch size while maintaining the same memory footprint when integrated with KIVI. Model weights and KV cache are the primary memory components accessed during generation. By effectively reducing the memory footprint of the KV cache, our method alleviates the memory bottleneck, enabling faster generation speeds as shown in Table 10.
Pruning Channels of Both Key and Value Cache. In this part, we explore the impact of pruning channels in the value cache (Appendix D). Specifically, for KV cache compression methods, we apply different pruning ratios to the channels of the key and value caches. Table 8 presents the results with LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2 on the LongBench benchmark. When evaluating the base model LLaMA-3-8B-Instruct, H2O or SnapKV with both key and value channel pruning perform comparably to H2O or SnapKV without pruning. In certain instances, the models with key and value channel pruning even outperform their non-pruned counterparts. For the base model Mistral-7B-Instruct-v0.2, pruning the value cache channels leads to a slight performance drop. This aligns with the observations in Figure 4, where the Key cache shows highly unbalanced magnitudes along the channel dimension, while the Value cache exhibits more uniform magnitudes. This suggests that the Value cache has less channel sparsity, making it harder to identify redundant channels for pruning. Nevertheless, pruning Value cache channels still contributes to further memory reduction in the KV cache.
@@ 5 Related Work
KV Cache Eviction. StreamingLLM (Xiao et al., 2023b) retains a few initial tokens along with recent tokens based on the observation of attention sinks, which can leading to the loss of critical information carried by the dropped tokens. H2O (Zhang et al., 2024c) selectively retains a small subset of tokens by greedily dropping those with lower contributions to cumulative attention. SnapKV (Li et al., 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts. FastGen (Ge et al., 2023) adaptively evicts tokens from attention heads that focus on local contexts, discarding non-special tokens that surround key tokens, while standard KV cache is applied to attention heads that attend more broadly. PyramidKV (Zhang et al., 2024b) and PyramidInfer (Yang et al., 2024) take a hierarchical approach, adjusting KV cache sizes across different layers by allocating more cache to lower layers and less to higher ones.
@@
Structured Pruning of LLMs. Traditional structured pruning  (Ma et al., 2023; Ding et al., 2023; Lu et al., 2024a) of LLMs typically focuses on removing unimportant layers, heads, or hidden dimensions, often leading to significant performance degradation. In contrast, our approach preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache. By dynamically identifying unimportant channels using data dependant criterion, our method greatly reduce the key cache size with minimal performance loss.
6 Conclusion and Limitations
Motivated by the observation that certain channels exhibit significantly larger magnitudes and the low-rank nature of the key cache (as shown by singular value analysis), we propose ThinK as a query-dependent pruning method targeting key cache channels. Optimized based on attention scores, our approach preserves essential information for each input query. ThinK seamlessly integrates with existing token-level KV cache pruning (Li et al., 2024; Zhang et al., 2024c) and KV cache quantization (Liu et al., 2024b), further improving inference efficiency. Extensive experiments on LongBench and Needle-in-a-Haystack benchmarks demonstrate its effectiveness, achieving comparable or superior performance to baselines while reducing key cache size by 40%. However, our method slightly increases TTFT (Time-to-First-Token) due to channel importance computation (see Appendix F). Additionally, while effective for key cache pruning, value cache pruning remains an open direction for future work.